A market consists of four individuals A, B, C and D whose demand functions are: -




QA=70-2P




QB=200-4P




QC=20-0.5P




QD=240-4P




If the industry supply equation is given by ๐‘„5 =40+3.5๐‘ƒ




Compute the equilibrium price and quantity and determine the quantities that will be



purchased by each individual.

1
Expert's answer
2022-02-14T15:18:41-0500

Total quantity demanded"=Qa+Qb+Qc+Qd"

"=70-2p+200-4p+20-0.5p+240-4p"

"=530-10.5p"

At equilibrium, total quantity demanded is equal to quantity supplied.

"530-10.5p=40+3.5p"

"530-40=3.5p+10.5p"

"490=14p"

"P=35"

"Q=530-10.5\u00d735=162.5"

Quantity consumed by each individual:

"Qa=70-2p=70-2\u00d735=0Units"

"Qb=200-4p=200-4\u00d735=130Units"

"Qc=20-0.5p=20-0.5\u00d735=2.5Units"

"Qd=240-4p=240-2\u00d735=170units"